OpenStudy (anonymous):

Complete the identity. Sin(α+β )/cos α cosβ = ? A. cot α + cot β B. tan α + tan β C. tan β + tan α D. -tan α + cot β

OpenStudy (anonymous):

write down the formula for \[\sin(\alpha+\beta)\] then divide

OpenStudy (anonymous):

@studystagram

OpenStudy (anonymous):

@satellite73 can you help me work out the problem all together show me how to do it

OpenStudy (studystagram):

B is the ans

OpenStudy (anonymous):

Thank you

OpenStudy (studystagram):

sin(A+B)/cosAcosB =(SinACosB + CosASinB) / CosACosB = SinACosB/CosACosB + CosASinB/ CosACosB = SinA/CosA + SinB/CosB sin(A+B)/cosAcosB = TanA + TanB

OpenStudy (anonymous):

Thank you so much!